# Spineless Traversal for Layout Invalidation

### v5

@@ 1. Introduction
Browsers use the “Double Dirty Bit” algorithm to find dirty nodes more quickly (Panchekha and Harrelson, 2024; Garseil, [n. d.]). This algorithm adds summary bits to identify and skip subtrees without any dirty elements. While this reduces the search time, Double Dirty Bit still has to traverse the tree, starting from the root, to find dirty elements, and thus accesses not only the actually-dirty elements but many extra “auxiliary nodes”. On large pages, auxiliary nodes can significantly outnumber the actually-dirty elements; and since each node access can cause a cache miss, simply traversing these auxiliary nodes, just to check their summary bits, can stall the layout algorithm for hundreds of microseconds. This problem is widely observed in practice; Google’s widely-used web performance tool, Lighthouse, measures tree depth and maximum children count precisely because these parameters determine the number of auxiliary nodes.
@@ 2.2. Formal Modeling
Luckily, the programming languages community has developed a sophisticated understanding of layout. Early work by Meyerovich and others (Jones et al., [n. d.]; Meyerovich and Bodik, 2010; Meyerovich, 2013) developed “attribute grammars” as a formalism for defining layout rules and implementations. The later Cassius project (Panchekha and Torlak, 2016; Panchekha et al., 2018; Prazina et al., 2023) developed a full, standards-compliant implementation of a significant fragment of web layout in this formalism, and the Hecate and Medea tools (Chen et al., 2022; Liu et al., 2023) have shown that automatic synthesis can be scaled to such fragments.
@@
A key property of attribute grammar DSLs like this one is that dependencies and traversal orders are static. Specifically, for any field assignment, examining the expression $T$ reveals which other fields on which other nodes it depends on. These static dependencies are critical for invalidation. Also, the schedule language ensures that the relative order in which two fields are computed never changes, even as nodes are added or removed. This is critical for Spineless Traversal. In this paper, we assume that the rules have no cyclic dependencies and that the schedule respects field dependencies; we also assume that the schedule has already been optimized by fusing traversals and ordering field assignments. A substantial literature exists on these topics (grafter; Chen et al., 2022; Liu et al., 2023); the layout implementation in our evaluation is detailed in Section 7.1. In any case, while this paper focuses on web layout, we expect Spineless Traversal to be applicable to incremental computations in other domains, including in compilation, static analysis, computer graphics, and databases, as long as they can be expressed as an attribute grammar in this or a similar DSL.
@@ 3.3. The Double Dirty Bit Algorithm
Skipping subtrees with no dirty bits reduces the number of auxiliary accesses and greatly improves performance (Garseil, [n. d.]; Panchekha and Harrelson, 2024). However, some auxiliary accesses remain. Specifically, when a dirty bit is set on some node, summary bits will be set for all ancestors of that node, which Double Dirty Bit will have to check. Moreover, for each ancestor, Double Dirty Bit will recurse into its children, adding even more auxiliary accesses. Since layout trees are poorly balanced and often very wide, the number of auxiliary accesses can be large. For example, a single dirtied node in Figure 1 has 66 auxiliary nodes, and on larger pages the dirty-to-auxiliary ratio can be even worse.
Developers can sometimes reduce the number of auxiliary accesses by reorganizing large, complex pages; performance monitoring tools like Google’s Lighthouse (Team, [n. d.]) will suggest doing so. But this can require global changes to the shape of the layout tree, which (in modern frameworks like React) requires refactoring the application as a whole. Naturally, an invalidation algorithm that simply did not require so many auxiliary nodes would be a superior solution. And while this is a quite specialized performance problem, it is a major source of latency in existing web browsers, which in turn are both critical application platforms and also already highly optimized, meaning remaining sources of latency are particularly challenging.
@@ 4.3. Order Maintenance
The final challenge is efficiently assigning timestamps to every dirty bit on every node. Simple incrementing integer timestamps, sadly, don’t work: inserting a node would requiring adjusting all later timestamps, which would introduce its own set of auxiliary accesses. Instead, following SAC (Acar, 2005), Spineless Traversal uses an order maintenance data structure (OM) to assign timestamps. First introduced by Dietz (1982), order maintenance maintains a totally ordered set of objects while allowing objects to be added and removed from the order arbitrarily. Crucially, adding, removing, and comparing nodes takes $O(1)$ time. Abstractly, order maintenance provides the following API:
@@ 5.3. Code Generation
Megatron’s code generation has three key steps: generating the node data structure, generating the field value computations, and generating the top-level recompute function. To simplify code generation and optimization, the compiler is organized as a staged interpreter (MetaOCaml). That is, we first implemented an interpreter, and then added staging annotations that generate a custom IR; Figure 10 summarizes the approach. The custom IR is then converted to C++ using traditional compiler optimizations. This approach was critical, as we found compiler correctness challenging: flaws in the dependency analysis could have unintuitive, long-distance consequences on program behavior that were difficult to distinguish from incorrect optimizations. The staging step allowed for much easier debugging.
@@ 6.3. Branchless Order Maintenance Comparison
Priority queue pushes and pops spend basically all of their time comparing order maintenance objects. Order maintenance objects are small and, thanks to our allocator, typically in cache. However, order maintenance object comparison has two cases (same or different second-level lists) and the pipeline stall from the conditional ends up being a bottleneck. Moreover, the branch predictor does not help much, because (thanks to the priority queue) the comparison is unpredictable. We therefore implemented a branchless comparison function, relying on inline assembly cmov instructions, shown in Figure 12; it executes in about five cycles of latency. Assembly makes the implementation non-portable, but this function is critical for performance and we weren’t able to make the C++ compiler generate comparably-fast code, despite several attempts.
@@ 8. Related Work
Speeding up computations by reusing previously-computed results is a long-studied topic in computer science broadly (Michie, 1968) and programming language theory in particular; Liu (2024) and Ramalingam and Reps (1993) give thorough surveys of the field. The recent Self-Adjusting Computation (SAC) (Acar, 2005) framework proposes incrementalizing arbitrary computations, including a cost semantics (Acar, 2009), optimizations for data structure operations (Acar et al., 2010), and opportunities for parallelization (Anderson et al., 2021). The Adapton framework (Hammer et al., 2014) aims at demand-driven incremental computation, and allows manually-specified annotations (Hammer et al., 2015) for greater reuse. While this prior work focuses on general-purpose computations, Spineless Traversal is focused on a particularly critical application: web browser layout. This application-specific focus has precedent: Pugh and Teitelbaum (1989) speed up memoization for functional programs over lists using “chunky decomposition”, while differential dataflow (McSherry et al., 2013), which incrementalizes relational algebra in databases, is prominent in industry.
Reps (1982) wrote the earliest work on incremental evaluation of attribute grammars, motivated by syntax-directed editors; later work (Reps et al., 1986) allows references to non-neighbor attributes. However, these early papers require recomputation immediately after every tree change, whereas web browsers, our target application, batch multiple updates and perform layout only once per frame. The standard in web browsers is instead the Double Dirty Bit algorithm, described in industry publications (Garseil, [n. d.]) and textbooks (Panchekha and Harrelson, 2024).
The formal methods community has put significant effort into formalizing web page layout. Jones et al. ([n. d.]) proposed using attribute grammars, similar to the DSL in Figure 3, for formalizing web-like layout rules. Later work (Chen et al., 2022) proposes synthesizing schedules from the attribute grammar rules, including proposals (Meyerovich and Bodik, 2010; Meyerovich, 2013) to use parallel schedules to further improve layout performance, though these proposals have not proven practical (servo-no-parallel). The Cassius project (Panchekha and Torlak, 2016) formalizes a significant fragment of CSS 2.1 using an attribute-grammar-like formalism. Our layout implementation is based on Cassius. Later work also proposed using the Cassius formalism to verify web page layouts (Panchekha et al., 2018), including in a custom proof assistant (Prazina et al., 2023). However, none of these works investigate incremental layout. By contrast, the Medea project (Liu et al., 2023) proposed synthesizing incremental layout algorithms by automatically synthesizing dirty bit propagation code. Our work extends Medea by exploring optimized incremental traversal algorithms.

### v7

@@ 1. Introduction
Browsers use the “Double Dirty Bit” algorithm to find dirty nodes more quickly (Panchekha and Harrelson, 2025; Garseil, 2009). This algorithm adds summary bits to identify and skip subtrees without any dirty elements. While this reduces the search time, Double Dirty Bit still has to traverse the tree, starting from the root, to find dirty elements, and thus accesses not only the actually-dirty elements but many extra “auxiliary nodes”. On large pages, auxiliary nodes can significantly outnumber the actually-dirty elements; and since each node access can cause a cache miss, simply traversing these auxiliary nodes, just to check their summary bits, can stall the layout algorithm for hundreds of microseconds. This problem is widely observed in practice; Google’s widely-used web performance tool, Lighthouse, measures tree depth and maximum children count precisely because these parameters determine the number of auxiliary nodes.
@@ 2.2. Formal Modeling
Luckily, the programming languages community has developed a sophisticated understanding of layout. Early work by Meyerovich and others (Jones et al., 2009; Meyerovich and Bodik, 2010; Meyerovich, 2013) developed “attribute grammars” as a formalism for defining layout rules and implementations. The later Cassius project (Panchekha and Torlak, 2016; Panchekha et al., 2018; Prazina et al., 2023) developed a full, standards-compliant implementation of a significant fragment of web layout in this formalism, and the Hecate and Medea tools (Chen et al., 2022; Liu et al., 2023) have shown that automatic synthesis can be scaled to such fragments.
@@
A key property of attribute grammar DSLs like this one is that dependencies and traversal orders are static. Specifically, for any field assignment, examining the expression $T$ reveals which other fields on which other nodes it depends on. These static dependencies are critical for invalidation. Also, the schedule language ensures that the relative order in which two fields are computed never changes, even as nodes are added or removed. This is critical for Spineless Traversal. In this paper, we assume that the rules have no cyclic dependencies and that the schedule respects field dependencies; we also assume that the schedule has already been optimized by fusing traversals and ordering field assignments. A substantial literature exists on these topics (Sakka et al., 2019; Chen et al., 2022; Liu et al., 2023); the layout implementation in our evaluation is detailed in Section 7.1. In any case, while this paper focuses on web layout, we expect Spineless Traversal to be applicable to incremental computations in other domains, including in compilation, static analysis, computer graphics, and databases, as long as they can be expressed as an attribute grammar in this or a similar DSL.
@@ 3.3. The Double Dirty Bit Algorithm
Skipping subtrees with no dirty bits reduces the number of auxiliary accesses and greatly improves performance (Garseil, 2009; Panchekha and Harrelson, 2025). However, some auxiliary accesses remain. Specifically, when a dirty bit is set on some node, summary bits will be set for all ancestors of that node, which Double Dirty Bit will have to check. Moreover, for each ancestor, Double Dirty Bit will recurse into its children, adding even more auxiliary accesses. Since layout trees are poorly balanced and often very wide, the number of auxiliary accesses can be large. For example, a single dirtied node in Figure 1 has 66 auxiliary nodes, and on larger pages the dirty-to-auxiliary ratio can be even worse.
Developers can sometimes reduce the number of auxiliary accesses by reorganizing large, complex pages; performance monitoring tools like Google’s Lighthouse (Team, 2024) will suggest doing so. But this can require global changes to the shape of the layout tree, which (in modern frameworks like React) requires refactoring the application as a whole. Naturally, an invalidation algorithm that simply did not require so many auxiliary nodes would be a superior solution. And while this is a quite specialized performance problem, it is a major source of latency in existing web browsers, which in turn are both critical application platforms and also already highly optimized, meaning remaining sources of latency are particularly challenging.
@@ 4.3. Order Maintenance
The final challenge is efficiently assigning timestamps to every dirty bit on every node. Simple incrementing integer timestamps, sadly, don’t work: inserting a node would requiring adjusting all later timestamps, which would introduce its own set of auxiliary accesses. Instead, following SAC (Acar et al., 2005), Spineless Traversal uses an order maintenance data structure (OM) to assign timestamps. First introduced by Dietz (1982), order maintenance maintains a totally ordered set of objects while allowing objects to be added and removed from the order arbitrarily. Crucially, adding, removing, and comparing nodes takes $O(1)$ time. Abstractly, order maintenance provides the following API:
@@ 5.3. Code Generation
Megatron’s code generation has three key steps: generating the node data structure, generating the field value computations, and generating the top-level recompute function. To simplify code generation and optimization, the compiler is organized as a staged interpreter (Kiselyov, 2018). That is, we first implemented an interpreter, and then added staging annotations that generate a custom IR; Figure 10 summarizes the approach. The custom IR is then converted to C++ using traditional compiler optimizations. This approach was critical, as we found compiler correctness challenging: flaws in the dependency analysis could have unintuitive, long-distance consequences on program behavior that were difficult to distinguish from incorrect optimizations. The staging step allowed for much easier debugging.
@@ 6.3. Branchless Order Maintenance Comparison
By using a binary heap, which has great cache locality, priority queue pushes and pops spend basically all of their time comparing order maintenance objects. Order maintenance objects are small and, thanks to our allocator, typically in cache. However, order maintenance object comparison has two cases (same or different second-level lists) and the pipeline stall from the conditional ends up being a bottleneck. Moreover, the branch predictor does not help much, because (thanks to the priority queue) the comparison is unpredictable. We therefore implemented a branchless comparison function, relying on inline assembly cmov instructions, shown in Figure 12; it executes in about five cycles of latency. Assembly makes the implementation non-portable, but this function is critical for performance and we weren’t able to make the C++ compiler generate comparably-fast code, despite several attempts. If portability is a concern, we also provided a branchless version implemented using bit operations.
@@ 8. Related Work
Speeding up computations by reusing previously-computed results is a long-studied topic in computer science broadly (Michie, 1968) and programming language theory in particular; Liu (2024) and Ramalingam and Reps (1993) give thorough surveys of the field. The recent Self-Adjusting Computation (SAC) (Acar et al., 2005) framework proposes incrementalizing arbitrary computations, including a cost semantics (Acar, 2009), optimizations for data structure operations (Acar et al., 2010), and opportunities for parallelization (Anderson et al., 2021). The Adapton framework (Hammer et al., 2014) aims at demand-driven incremental computation, and allows manually-specified annotations (Hammer et al., 2015) for greater reuse. While this prior work focuses on general-purpose computations, Spineless Traversal is focused on a particularly critical application: web browser layout. This application-specific focus has precedent: Pugh and Teitelbaum (1989) speed up memoization for functional programs over lists using “chunky decomposition”, while differential dataflow (McSherry et al., 2013), which incrementalizes relational algebra in databases, is prominent in industry.
Reps (1982) wrote the earliest work on incremental evaluation of attribute grammars, motivated by syntax-directed editors; later work (Reps et al., 1986) allows references to non-neighbor attributes. However, these early papers require recomputation immediately after every tree change, whereas web browsers, our target application, batch multiple updates and perform layout only once per frame. The standard in web browsers is instead the Double Dirty Bit algorithm, described in industry publications (Garseil, 2009) and textbooks (Panchekha and Harrelson, 2025).
The formal methods community has put significant effort into formalizing web page layout. Jones et al. (2009) proposed using attribute grammars, similar to the DSL in Figure 3, for formalizing web-like layout rules. Later work (Chen et al., 2022) proposes synthesizing schedules from the attribute grammar rules, including proposals (Meyerovich and Bodik, 2010; Meyerovich, 2013) to use parallel schedules to further improve layout performance, though these proposals have not proven practical (Project, 2023). The Cassius project (Panchekha and Torlak, 2016) formalizes a significant fragment of CSS 2.1 using an attribute-grammar-like formalism. Our layout implementation is based on Cassius. Later work also proposed using the Cassius formalism to verify web page layouts (Panchekha et al., 2018), including in a custom proof assistant (Prazina et al., 2023). However, none of these works investigate incremental layout. By contrast, the Medea project (Liu et al., 2023) proposed synthesizing incremental layout algorithms by automatically synthesizing dirty bit propagation code. Our work extends Medea by exploring optimized incremental traversal algorithms.